# A Fourier-based inference method for learning interaction kernels in particle systems

### v1

@@ 4.2 Inference of interaction kernels
[…]
@@ 5 Conclusion
    Our approach is easy to implement, since it only requires the approximation of the moments of the invariant measure and the solution of a low-dimensional linear system, and computationally cheap. Moreover, because of its versatility, it can infer complex interaction kernels. On the other hand, its main limitation, illustrated in the convergence analysis, is the dependence of the approximation error on the number $K$ of Fourier coefficients used in the expansion of the interaction kernel. In fact, larger values of $K$, which in principle should provide better estimates, can potentially lead to worse results if the observed trajectory or the number of particles in the system are not sufficiently large.
@@
    The work presented in this paper can be extended in several directions. First, we would like to improve the convergence result by quantifying the dependence on the number $K$ of Fourier coefficients, as this could help in finding techniques to reduce the approximation error. Moreover, we believe that it should be possible to define large classes of diffusion processes for which Assumption 2.4 holds and $\delta(K)\to 0$ in Theorem 3.8, such that the asymptotic unbiasedness of the estimator is guaranteed a priori by a theoretical result. Another interesting development would be lifting the regularity hypotheses in Assumption 3.1 on the confining and interaction potentials and considering the less regular functions to be estimated. In this paper, we consider the problem in one dimension in space. Similarly to the method of moments studied in [40], we expect that the methodology developed in this paper could be extended to the multidimensional case, considering the appropriate tensor products of orthogonal polynomials with respect to the invariant measure. The detailed analysis of the multidimensional problem will be presented elsewhere. Finally, the numerical experiments suggest that Assumption 1.1 on the uniqueness of the invariant measure of the mean-field dynamics, e.g. our assumption that we have uniform propagation of chaos, is not necessary and that our semiparametric method is applicable even in the presence of multiple stationary states. This was also demonstrated numerically for the eigenfunction martingale estimator in [39]. We expect that the results from [34], combined with the linearization approach from [41] are sufficient to provide a rigorous justification of our methodology to the case where the mean-field dynamics exhibits phase transitions. Finally, it would be interesting to apply our method to the kinetic Langevin / hyperelliptic setting [4, 26, 25]. All these interesting problems will be studied in future work.

### v2

@@ 4.3 Inference of interaction kernels
    All of these interaction potentials give rise to multiple stationary states at low temperatures. We note that $W_{3}$ is somewhat similar to the Morse potential, where we replaced the radius with $x^{2}$, and it is still the difference between an attractive and a repulsive potential, and $W_{4}$ represents a Gaussian attractive interaction. Moreover, we set $D=5$, $a=0.5$, $r=1$ in $W_{3}$, and $A=5$ in $W_{4}$. We consider $N=250$ particles and set the final observation time at $T=5000$ for the first two kernels, while we choose $N=500$ and T=10  [^0] T=10\,000 for $W_{3}$ and $W_{4}$. We then compute the estimators $(\widehat{W}^{\prime}_{i})_{T,N}^{(K)}$ for all $i=1,\dots,4$, and report the results in Figure˜6. In the first test case, where the interaction kernel $W_{1}^{\prime}$ is polynomial, the approximation obtained with the Fourier coefficients $K=5$ is accurate, especially in the interval [2,2][-2,2] where there are more observations. Similar considerations can be made for the second test case, where we still use $K=5$ Fourier coefficients, even if the function $W_{2}^{\prime}$ to be inferred is not polynomial. The third and fourth examples are more challenging, as they behave differently at infinity. In fact, both $W_{3}^{\prime}$ and $W_{4}^{\prime}$ vanish for $x\to\pm\infty$, and therefore they cannot be well approximated by any polynomial of finite degree. We observe that even if the estimators, which are computed with the $K=11$ and $K=9$ Fourier coefficients, respectively, are not as accurate as in the previous test cases, in particular for larger values of $x$, they are still able to match the overall shape of the interaction kernels. These numerical experiments show the potentiality of the approach presented in this work, which allows us to obtain a reasonable reconstruction of the interaction kernel using only a finite single trajectory from the interacting particle system, even for more complex scenarios that do not fit in the theoretical analysis.
@@ 5 Conclusion
    Our approach is easy to implement, since it only requires the approximation of the moments of the invariant measure and the solution of a low-dimensional linear system, and computationally cheap. Moreover, because of its versatility, it can infer complex interaction kernels. On the other hand, its main limitation, illustrated in the convergence analysis, is the dependence of the approximation error on the number $K$ of Fourier coefficients used in the expansion of the interaction kernel. In fact, larger values of $K$, which in principle should provide better estimates, can potentially lead to worse results if the observed trajectory or the number of particles in the system are not sufficiently large. The optimal choice of $K$ depends on multiple interacting factors, including the regularity of the function $W^{\prime}$ to be approximated, the observation time $T$, the number of particles $N$, and numerical aspects such as the conditioning of the associated linear system and the stability of the Gram–Schmidt procedure. Therefore, developing and testing adaptive, data-driven procedures for automatically selecting $K$, such as the criterion inspired by the classical bias–variance trade-off in [14, Section 5], could significantly improve the robustness and practical applicability of our methodology.
@@
    The work presented in this paper can be extended in several other directions. First, we would like to improve the convergence result by quantifying the dependence on the number $K$ of Fourier coefficients, as this could help in finding techniques to reduce the approximation error. Moreover, we believe that it should be possible to define large classes of diffusion processes for which Assumption˜2.4 holds and $\delta(K)\to 0$ in Theorem˜3.10, such that the asymptotic unbiasedness of the estimator is guaranteed a priori by a theoretical result. Another interesting development would be lifting the regularity hypotheses in Assumption˜3.1 on the confining and interaction potentials and considering the less regular functions to be estimated. In this paper, we consider the problem in one dimension in space. Similarly to the method of moments studied in [45], we expect that the methodology developed in this paper could be extended to the multidimensional case. Conceptually, we believe that the framework of representing the interaction kernel in terms of an orthogonal basis with respect to the invariant measure of the mean-field dynamics remains applicable. Moreover, the construction of the linear system for computing the Fourier coefficients of the expansion could also be extended relatively straightforwardly. Nevertheless, moving to high-dimensional particle systems introduces, in addition to a more complicated theoretical analysis, several technical challenges that make this extension nontrivial. First, the construction of the orthogonal polynomial basis with respect to the invariant measure, which plays a central role in our estimator, becomes significantly more involved in higher dimensions. In particular, the initial basis of monomials would need to be replaced by tensor products of monomials, whose number grows exponentially with the dimension due to the curse of dimensionality. This leads to a more elaborate orthogonalization procedure, where the Gram–Schmidt algorithm might need to be replaced by more stable methods such as QR factorization with Householder reflections. Moreover, the increase in basis dimension implies larger linear systems to solve for the Fourier coefficients, which in turn affects both the conditioning of the system and the computational cost. The detailed analysis of the multidimensional problem will be presented elsewhere. To move toward more realistic applications, we will need consider (trajectory) observations that are contaminated by noise. This could be done, in principle, by combining our inference methodology with techniques from filtering, or by setting up the inference problem as a Bayesian inverse problem, as done in [40]. We expect that incorporating observation noise will require additional steps in the inference methodology, increase the complexity of the statistical analysis, and potentially affect convergence rates. Finally, the numerical experiments suggest that Assumption˜1.1 on the uniqueness of the invariant measure of the mean-field dynamics, e.g. our assumption that we have uniform propagation of chaos, is not necessary and that our semiparametric method is applicable even in the presence of multiple stationary states. This was also demonstrated numerically for the eigenfunction martingale estimator in [44]. We expect that the results from [38], combined with the linearization approach from [46] are sufficient to provide a rigorous justification of our methodology to the case where the mean-field dynamics exhibits non-uniqueness of stationary states. However, we expect that the performance of our estimator deteriorates as we approach the critical noise strength where the phase transition occurs. In fact, as shown in [38], at the critical point the convergence to the invariant measure is only algebraic and the empirical time averages may converge more slowly. Finally, it would be interesting to apply our method to the kinetic Langevin / hypoelliptic setting [4, 27, 26]. All these interesting problems will be studied in future work.